OpenStudy (anonymous):

What is the equation, in standard form of a parabola that contains the following points. (-2,24) (0,4) (4,60)

OpenStudy (tkhunny):

f(x) = ax^2 + bx + c Create these: f(-2) = 24 f(0) = 4 f(4) = 60 f(0) = 4 give x = 4 f(x) = ax^2 + bx + 4 Create these: f(-2) = 24 f(4) = 60 And solve the small system for a and b.